18. Solve the system of equations shown below. Use any method you wish. 2x2 - xy + y2 = 8 xy = 4 {2x?
Answer #1

Given (Ф 2 х”— 2y ty“. ху = 9 п Х. ч У x= 4 in equation © substitute + у — 8 2() - (9) +y. 8 2 (1) - 32 – чу” + y1. y? 32 -y - 4 y & y = + Vs Ya Iru y=+2n y = 12 - ya -2, 2, -282, 282 2 Substitute y=-2 in equation 26 (-2) 24 2 -2 substitute y=2 inThe solution to the system is the complete set of ordered pains that pains that are valid solutions (-2,-2), (2,2), (-2,-2/2)